When a bank makes its loans, if it screens its borrowers it will collect repayment revenue of $4,000 per loan, but if it doesn't screen its borrowers then it will collect $4,000 per loan with probability 0.85 and collect $1,000 with probability 0.15. The cost of screening is $200 per loan. In this case, the expected payoff for the bank from screening is ____ and the expected payoff from not screening is ____

Group of answer choices

$3,200; $3,550

$3,200; $3,400

$3,800; $3,400

$3,800; $3,550
Answer #1

In case of screening, it will collect 4000-200 = 3800

In case of not screening, it will collect 4000*0.85 + 1000*0.15 = 3400+150 = 3550

option(D)